Rajat gave a quarter of his chocolates to Harsh and then gives 5 of his chocolates to Tushar.He has7 left.How many did he have to start with?











Answer by mananth(16946)

rajat had x chocolates
gives 1/4 to harsh
so he gives x/4 to harsh
..
he has balance x-(x/4)
=3x/4
gives 5 to tushar
(3x/4) -5 = 7
LCD =4
multiply equation by 4
3x-20=28
3x=20+28
3x=48
x=16